# OmniGS: Fast Radiance Field Reconstruction using Omnidirectional Gaussian Splatting

### v3

@@ Abstract
Photorealistic reconstruction relying on 3D Gaussian Splatting has shown promising potential in various domains. However, the current 3D Gaussian Splatting system only supports radiance field reconstruction using undistorted perspective images. In this paper, we present OmniGS, a novel omnidirectional Gaussian splatting system, to take advantage of omnidirectional images for fast radiance field reconstruction. Specifically, we conduct a theoretical analysis of spherical camera model derivatives in 3D Gaussian Splatting. According to the derivatives, we then implement a new GPU-accelerated omnidirectional rasterizer that directly splats 3D Gaussians onto the equirectangular screen space for omnidirectional image rendering. We realize differentiable optimization of the omnidirectional radiance field without the requirement of cube-map rectification or tangent-plane approximation. Extensive experiments conducted in egocentric and roaming scenarios demonstrate that our method achieves state-of-the-art reconstruction quality and high rendering speed using omnidirectional images. The code will be publicly available.
@@ 1 Introduction
Recently, 3D Gaussian Splatting (3DGS) [11] effectively addresses the limitation of NeRF by introducing 3D Gaussians to explicitly represent radiance field representation. Each 3D Gaussian is a point associated with certain attributes, i.e. position, color, scale, rotation, and opacity. For rendering, the EWA splatting algorithm [23] is applied to project and rasterize 3D Gaussians onto the screen space. Benefiting from the highly efficient rendering process, 3DGS significantly reduces training and inference time in high-quality radiance field reconstruction, having great potential in real-time applications [10, 21, 22, 7]. Nevertheless, the current splatting algorithm is only compatible with undistorted perspective image rendering.
@@ 3.4 Backward Optimization
where $M$ is the total number of instances generated by this Gaussian in all tiles, $\mathbf{R}$ is the rotation matrix converted from quaternion $\mathbf{q}$. Vectors $\mathbf{s}$ and $\mathbf{p}$ are the screen-space and image-space coordinates, respectively. We retain the common portion from gradients given by [11], and replace the following parts with our omnidirectional gradients: ${\frac{\partial\tilde{\mathbf{\Sigma}}}{\partial\mathbf{\Sigma}}},\displaystyle{\frac{\partial\mathbf{J}}{\partial\mathbf{t}}},\displaystyle{\frac{\partial\mathbf{p}}{\partial\mathbf{s}}},\displaystyle{\frac{\partial\mathbf{s}}{\partial\mathbf{t}}}$. More detailed derivation can be found in our supplementary material.
@@ 5.1 Implementation and Experiment Setup
We accomplished OmniGS based on LibTorch framework, which is the C++ version of PyTorch. The tile-based omnidirectional rasterizer was implemented with custom CUDA kernels. As for the usage of datasets, we directly utilized the calibrated camera poses and sparse SfM point clouds contained in the datasets as initial input. But notably, we performed openMVG [15] SfM on the OmniBlender scenes of EgoNeRF dataset, since these scenes provide no sparse point clouds. For the sake of a fair comparison, we used an RTX-3090 GPU to conduct all experiments on OmniGS, using the provided training and testing split, and gathered the RTX-3090 baseline results reported by the authors of datasets [6] and [5] unless specifically stated. We evaluated the results in terms of PSNR, SSIM, LPIPS, and rendering FPS, which are common criteria for photorealistic reconstruction. The baselines, except for EgoNeRF, were evaluated in terms of perspective novel-view synthesis, regarding their original objective. Detailed hyperparameter setups are discussed in our supplementary material.
@@ 5.4 Additional Perspective Rendering Evaluation
As shown in Tab. 3, when considering the tested-as-trained performance, i.e. the P-trained P-tested 3DGS and O-trained O-tested OmniGS, 3DGS was a little better than our OmniGS (P denotes perspective and O denotes omnidirectional). This is because when scene models are evaluated in the form of equirectangular images, the near-pole distortion of real omnidirectional cameras will deteriorate the quantitative results[5]. The above phenomenon indicates that we should unify the form of testing images before comparing OmniGS with the perspective 3DGS. Taking this into account, we re-evaluated the OmniGS results by cropping the rendered testing views into perspective images, and then compared the two P-tested results. It came out that our method was equipped with the ability to generate better perspective views cropped from the rendered omnidirectional images (PSNR$+$1.812, SSIM$+$0.012, LPIPS--0.045). In addition, we rendered the 3DGS-trained models with our omnidirectional rasterizer, then compared the two O-tested results. The omnidirectional novel-view synthesis quality of our models was also higher than that of 3DGS (PSNR$+$2.801, SSIM$+$0.055, LPIPS--0.103). Moreover, qualitative results (Fig. 6) figure out that the perspective 3DGS tended to lose detail and suffered artifacts due to the low utilization rate of observations. It can use only one limited view at one time. In contrast, when given the same time for training, OmniGS made use of the whole omnidirectional environment in each iteration, reaching more robust densification and faster model convergence. At the same time, the omnidirectional images rendered from OmniGS-reconstructed models can be cropped to generate perspective views properly, showing the strong scalability of our method.
@@ 6 Conclusion
Table 3: Quantitative comparison of perspective rendering results on 360Roam Dataset. P denotes perspective image, and O denotes omnidirectional image.

### v5

@@ Abstract
Photorealistic reconstruction relying on 3D Gaussian Splatting has shown promising potential in various domains. However, the current 3D Gaussian Splatting system only supports radiance field reconstruction using undistorted perspective images. In this paper, we present OmniGS, a novel omnidirectional Gaussian splatting system, to take advantage of omnidirectional images for fast radiance field reconstruction. Specifically, we conduct a theoretical analysis of spherical camera model derivatives in 3D Gaussian Splatting. According to the derivatives, we then implement a new GPU-accelerated omnidirectional rasterizer that directly splats 3D Gaussians onto the equirectangular screen space for omnidirectional image rendering. We realize differentiable optimization of the omnidirectional radiance field without the requirement of cube-map rectification or tangent-plane approximation. Extensive experiments conducted in egocentric and roaming scenarios demonstrate that our method achieves state-of-the-art reconstruction quality and high rendering speed using omnidirectional images. The code will be publicly available at https://github.com/liquorleaf/OmniGS.
@@ 1 Introduction
Recently, 3D Gaussian Splatting (3DGS) [11] effectively addresses the limitation of NeRF by introducing 3D Gaussians to explicitly represent radiance field. Each 3D Gaussian is a point associated with certain attributes, i.e. position, color, scale, rotation, and opacity. For rendering, the elliptical weighted average (EWA) splatting algorithm [23] is applied to project and rasterize 3D Gaussians onto the screen space. Benefiting from the highly efficient rendering process, 3DGS significantly reduces training and inference time in high-quality radiance field reconstruction, having great potential in real-time applications [10, 21, 22, 7]. Nevertheless, the current splatting algorithm is only compatible with undistorted perspective image rendering.
@@ 3.4 Backward Optimization
where $M$ is the total number of instances generated by this Gaussian in all tiles. Note that $\mathbf{R}$ is the rotation matrix converted from quaternion $\mathbf{q}$, which is an inner property determining the Gaussian covariance along with $\mathbf{S}$, and is different from the rotation part $\mathbf{W}$ of transform $\mathbf{T}_{\text{cw}}$. Vectors $\mathbf{s}$ and $\mathbf{p}$ are the screen-space and image-space coordinates, respectively. We retain the common portion from gradients given by [11], and replace the following parts with our omnidirectional gradients: ${\frac{\partial\tilde{\mathbf{\Sigma}}}{\partial\mathbf{\Sigma}}},\displaystyle{\frac{\partial\mathbf{J}}{\partial\mathbf{t}}},\displaystyle{\frac{\partial\mathbf{p}}{\partial\mathbf{s}}},\displaystyle{\frac{\partial\mathbf{s}}{\partial\mathbf{t}}}$. Detailed derivation can be found in our supplementary material.
@@ 5.1 Implementation and Experiment Setup
We accomplished OmniGS based on LibTorch framework, which is the C++ version of PyTorch. The tile-based omnidirectional rasterizer was implemented with custom CUDA kernels. As for the usage of datasets, we directly utilized the calibrated camera poses and sparse SfM point clouds contained in the datasets as initial input. But notably, we performed openMVG [15] SfM on the OmniBlender scenes of EgoNeRF dataset, since these scenes provide no sparse point clouds. For the sake of a fair comparison, we used an RTX-3090 GPU to conduct all experiments on OmniGS, using the provided training and testing split, and gathered the RTX-3090 baseline results reported by the authors of datasets [6] and [5] unless specifically stated. We evaluated the results in terms of PSNR, SSIM, LPIPS, and rendering FPS (forward pass inference speed), which are common criteria for photorealistic reconstruction. The baselines, except for EgoNeRF, were evaluated in terms of perspective novel-view synthesis, regarding their original objective. Detailed hyperparameter setups are discussed in our supplementary material.
@@ 5.4 Additional Perspective Rendering Evaluation
As shown in Tab. 3, when considering the tested-as-trained performance, i.e. the P-trained P-tested 3DGS and O-trained O-tested OmniGS, 3DGS was a little better than our OmniGS (P denotes perspective and O denotes omnidirectional). This is because when scene models are evaluated in the form of equirectangular images, the near-pole distortion of real omnidirectional cameras will deteriorate the quantitative results[5]. The above phenomenon indicates that we should unify the form of testing images before comparing OmniGS with the perspective 3DGS. Taking this into account, we re-evaluated the OmniGS results by cropping the rendered testing views into perspective images, and then compared the two P-tested results. It came out that our method was equipped with the ability to generate better perspective views cropped from the rendered omnidirectional images (PSNR$+$1.812, SSIM$+$0.012, LPIPS--0.045). In addition, we rendered the 3DGS-trained models with our omnidirectional rasterizer, then compared the two O-tested results. The omnidirectional novel-view synthesis quality of our models was also higher than that of 3DGS (PSNR$+$2.801, SSIM$+$0.055, LPIPS--0.103). Moreover, qualitative results (Fig. 6) figure out that the perspective 3DGS tended to lose detail and suffered artifacts due to the low utilization rate of observations. It can use only one limited view at one time. In contrast, when given the same time for training, OmniGS made use of the whole omnidirectional environment in each iteration, reaching more robust densification and faster model convergence. Omnidirectional images rendered from OmniGS-reconstructed models can also be cropped to generate perspective views properly, showing the strong scalability of our method.
@@ 6 Conclusion
Acknowledgements: This work was supported by the China National Key Research and Development Program under Grant 2022YFB3903804.
@@
Table 3: Quantitative comparison of perspective rendering results on 360Roam Dataset. P denotes perspective image, and O denotes omnidirectional image. Though OmniGS performs slightly worse when tested as trained, it outperforms 3DGS when the form of testing images are unified into O or P, respectively.